IODP Proceedings    Volume contents     Search


Site M00281

Expedition 313 Scientists2


Transit to Hole M0028A

Preparations for transit to Integrated Ocean Drilling Program (IODP) Hole M0028A commenced at 2100 h Universal Time Coordinated (UTC) on 22 May 2009. The jack-down procedure began at 2315 h, and the L/B Kayd moved off IODP Hole M0027A at 2335 h, arriving at Hole M0028A at 0100 h on 23 May. The legs were lowered to the seafloor by 0120 h, when the preloading procedure began. The drilling floor was opened up to operations personnel at 1100 h to begin setting up generators, powering up containers, and preparing the drilling floor.

Hole M0028A

The casing plan for Hole M0028A was to run the casing as deep in the hole as possible. The casing operation began after midday and continued overnight (Table T1). Progress was very slow because of the ground conditions. Just before 0600 h on 24 May 2009, the casing twisted off at the crossover sub 1–2 m above the seabed. A remotely operated vehicle (ROV) inspection confirmed that casing was protruding from the seabed and lengths of casing pipe were lying on the seabed, which would require a salvage operation to remove.

The decision was made to try to restart the hole at the same location. The new casing string was prepared and run into the seabed, and, rather than case to a great depth, it was decided to run in open hole using the PQ string to the target depth of 215 m drilling depth below seafloor (DSF) and then continuously core the middle and lower sections of the hole to 750 m DSF, as this was the primary objective. By the end of 24 May, 80 m had been drilled with the noncoring inner barrel inserted.

At 0030 h on 25 May, recovery of the noncoring barrel (to flush the pipe) began. However, the drill pipe became stuck at 102 m DSF. The jam was cleared by 0830 h, and open-hole drilling resumed, reaching ~157 m DSF by 1450 h. Drilling stopped at this point to lower the casing further into the seabed using the L/B Kayd before restarting. Just before 1800 h, a clay formation was encountered at 178 m DSF. The core barrel was inserted, but the drill pipe was sticking, so the base of the borehole was reamed out. This proved unsuccessful. Sticking points had been encountered higher up in the hole (probably due to swelling clays), so a much more substantial reaming operation began in order to condition the hole before the start of continuous coring operations. The pipe was tripped back to 98 m DSF, and a mud mix that would help inhibit the swelling clays was pumped into the hole.

The reaming operation continued overnight and for most of the following morning. The hole was open-hole drilled to 220 m DSF by 1230 h on 26 May, and coring began using the standard rotary corer (ALN) drill string. Backpressure was encountered during coring on the first and second runs. After that, backpressure diminished. A zone of high backpressure was encountered in Hole M0027A at a similar depth interval. Coring and reaming continued until midnight, with core recovery 100% or greater on all runs, reaching 246 m DSF. Overnight, core recovery became more variable as sandy formations were encountered, and there were problems with the drill pipe sticking, which required more hole conditioning. At 0820 h on 27 May at 264 m DSF, the core barrel did not latch in when lowered. It could not be retrieved, and there was no flush. The only solution was to trip the pipe and recover the barrel. The pipe was tripped by 1330 h, and the bottom of the bottom-hole assembly (BHA) was found to be blocked with sand. As the drill pipe had been tripped, this was an opportunity to try to deepen the casing beneath the seabed to improve its stability. The casing operation continued until midnight.

On 28 May, casing continued to be run, interrupted for a few hours by mechanical problems with the drilling rig. An ROV inspection at 0900 h showed that the casing had entered the seabed close to the previously abandoned casing and clear of any casing pipe debris. The operation was completed at 1945 h, by which time the foot of the casing was at 22.39 m DSF and occupying the hole from which the PQ pipe and casing had been previously tripped. The operation to lower the PQ drill pipe back to 264 m DSF and to clean and ream the hole continued overnight and was completed by 1945 h on 29 May.

The first cores of this new run were very sandy with poor recovery. There were also problems getting the core barrel to latch in because of sand entering the BHA. This required removal of some pipes to raise the bit from the bottom of the hole and pumping mud to clear the BHA. Once flushed clear, the core barrel was lowered and latched in, and then the drill pipe was lowered to the base of the hole to begin coring again at 2310 h. Backpressure due to artesian water was also noted in this zone.

By midnight on 29 May, the base of the hole had advanced to 270 m DSF. Coring continued throughout 30 May with the ALN core barrel, and by midnight, the hole had advanced to 317.62 m DSF. There were some problems with recovery and blocked core barrels due mainly to alternating beds of lithified sandstones and unlithified silty sands and clays. Zones of high backpressure were also encountered at 274 and 305 m DSF, but there was no evidence for gas being the cause.

From ~0800 h on 31 May, recovery dropped significantly as a coarse, sandy formation was encountered. By 1200 h the bottom of the hole had advanced to 339 m DSF. At this time the drill pipe (probably the BHA) got stuck in the hole following a cave-in. After some effort, the drill pipe was freed, but it was apparent that there was a problem with the bit on the BHA. There was little further penetration, and on retrieval, the inner barrel showed signs of damage and scoring. The drill pipe was tripped, and the BHA was on deck at 1740 h. The bit and reaming shell were missing, having probably been sheared off while freeing the drill pipe. Rather than attempt to restart drilling immediately with a new BHA, it was decided to fish for the missing parts, as they could impede further progress in the hole. The Bowen spear was made ready, and the fishing operation began at 2020 h.

The fishing operation proceeded very slowly throughout the night. At 0600 h on 1 June, the tool was at 198 m DSF and was encountering bridges that were difficult to wash away. Around 1200 h, the tool encountered a zone of swelling clays that was known to extend to 240 m DSF. Progress was halted by the clays, torque on the string was very high, and the drill string was not advancing. At 1230 h the decision was made to abandon the fishing operation, as there was little chance of further progress downhole and there was danger that the tool would also be lost. The tool was tripped and was back on deck at 1430 h. A new BHA was prepared with a diamond impregnated drill bit, capable of drilling through or pushing aside the missing bit and reaming shell. By midnight on 1 June, washing and reaming down had advanced the drill string to 143 m DSF, and by midnight on 2 June, it was at 326.77 m DSF.

Coring started at 0215 on 3 June and continued until midnight. The cored material was very variable for several core runs, from solid sandstone to running sand, with thin clay layers. The loose sand caused caving and slowed the progress for ~15 m, with several cores incorporating infill.

On 4 June, coring from 360 m DSF became easier, although core recovery was variable. Small cavings occasionally slowed progress through alternating soft and firm formations. From 1750 h onward, sandy formations forced shorter core runs and caving sands required additional reaming.

At ~0100 h on 5 June, a bridge formed above the BHA that caused the drill string to become stuck. Several hours were spent trying to pull the BHA past the bridge and circulate mud. This was unsuccessful, and the decision was made at 0900 h to continue coring using the HQ drill string through the PQ drill string, effectively using the PQ string as casing set at 404 m DSF. Before the HQ string was run into the hole, a through-pipe gamma ray wireline log was taken from 390 m DSF to the surface, which was completed by 1530 h. A total of 10 m of infill material in the PQ pipe had to be reamed out before HQ coring could begin.

At 0330 h on 6 June, coring using the HQ Tight Tolerance (HQTT) core barrel commenced. After four core runs, a hole collapse due to fine caving sands risked the HQ string becoming stuck, so the HQ string was pulled back into the PQ casing (0830–1320 h). Attempts were then made to ream back down, but the hole pressurized and four additional pipes had to be removed. A second attempt at reaming back down was successful, and coring started again at 1620 h. However, on the fifth HQTT core run, the drill string became stuck again for an hour. Amendments were subsequently made to the mud mixture.

The first few hours of 7 June were spent circulating and advancing the HQ string to the base of the PQ casing. While advancing the HQ string, the driller noted little penetration and thought there was an obstruction. The HQ string was pulled back to the surface and the BHA checked. The bit was still attached but was severely worn; it was replaced with a similar but harder matrix-impregnated bit. Upon running the HQ string down the hole, 67 m of sand was encountered inside the PQ string. The remainder of the day and the beginning of Monday, 8 June, was spent reaming out this material. A core of infill was taken from material believed to represent ~403.79–411.86 m DSF.

At 0745 h on 8 June, normal coring commenced, with nine good cores collected. The improvement in coring progress continued into Tuesday, 9 June. Coring was suspended from 1055 to 1225 h because of an electrical storm.

Coring continued to progress well until 1200 h on 10 June, with eight core runs and 18.5 m penetration, reaching 533.85 m DSF. However, a decreasing penetration rate caused concern, with the view that the inside cutting diameter of the core bit had worn out and thus the bit was cutting an oversized core. It was decided to pull the HQ string to check the bit. This decision coincided with the detection of H2S in the last core run returned to the deck (~1100 h). Precautions were taken, including venting the core prior to curation, and the borehole was monitored. The HQ string was pulled and the bit inspected. Although the bit showed considerable wear, it still did not account for the poor penetration. At 1725 h, the HQ string was run back into the hole with a polycrystalline diamond (PCD) bit attached. The hole appeared to be stable, and ~1.2 m of cavings was cleared. After reaming, the core barrel was retrieved for an H2S test (to 2400 h).

From 0035 to 0515 h on 11 June, there was a smell of H2S on the drill floor, although nothing registered on the gas monitors. The first core of the day smelled of H2S and registered 4.3 ppm on the gas monitor, which soon dissipated. Drilling continued with regular checks for H2S returning with the mud. The second and third cores also smelled of H2S but did not register on the gas monitor. Several more core runs were made, resulting in near 100% recovery, but failed to penetrate the full length of the 3 m core run because of a harder lithology. As the lithology changed to looser material, full core runs were recovered.

An electrical storm briefly halted operations from 0255 to 0330 h on 12 June. Penetration rates slowed in comparison to previous days, but with no discernible change in the lithology.

Early on 13 June, the drill string was pulled, and it was found that the outer cutting edge of the bit had worn away. At 0555 h, the HQ string was run back down with a new bit and stabilizer ring. At 326 m DSF, very high mud backpressure indicated a blocked bit. It was not possible to latch the core barrel with the overshot. A second HQ trip was made, and sand was found behind the core barrel, blocking its release. On running the HQ string back in, it was found that the PQ string had filled up with sand to 250 m DSF, and flushing was required. At ~439 m DSF, the drill string easily dropped into the hole.

The HQ string reached the base of the hole at 0530 h on 14 June, and coring recommenced with the HQTT barrel. Core recovery varied throughout the day but improved slightly on 15 June, when coring entered stiff, swelling clays.

At 0225 h on 16 June, the end of coring in Hole M0028A was declared (base of last core; 669 m DSF). At 0240 h, the final core was curated. The last 10 double rods (61 m) were lifted to ream out the bottom of the hole in readiness for logging. The wireline and vertical seismic profile (VSP) logging winches were set up on either side of the drill rig. At 1140 h, through-pipe spectral gamma ray logging began and was completed by 1650 h. A marine mammal watch began at 1730 h prior to the through-pipe VSP logging commencing at 1800 h.

The VSP tool was returned to the deck by 0030 h on 17 June. At 0130 h, mud circulation began, and the HQ string pulled to a clay layer ~30 m below the PQ string at 425 m DSF, keeping it below the running sand that caused earlier coring and reaming problems. Between 0450 and 2100 h, wireline logging and VSP of the lower open-hole interval from 669 to 425 m DSF was conducted with no problems. The air guns were left in the water, test firing until the next VSP run (through PQ pipe), which started at ~0200 h on 18 June. The marine mammal observation continued throughout this period.

At midnight on 18 June, the HQ string was pulled slowly out of Hole M0028A. The cautious approach was to prevent the hole from backfilling with sand, and this was confirmed when logging tools passed safely to the base of the PQ string. After successfully completing the VSP run at 0805 h, resistivity and acoustic sondes were run in the open-hole interval directly below the PQ string (396–425 m DSF). Various attempts were then made to lift the PQ string, which was found to be stuck fast and likely to be held higher up in the hole. Attempts were made to cut the PQ pipe at 387 and 351 m DSF using a cutting tool, which finally succeeded at 286 m DSF before the HQ string and cutter also became stuck.

The HQ rods were worked through the night into 19 June in an effort to free the string. Meanwhile, a spare cutter was ordered from the shore. Other methods to free the HQ string were explored: a heavy tool was manufactured that was intended to break the tool joint connecting the cutter to the string, and a pinch bar was attached to the wireline and repeatedly hammered, through freefall, onto the top of the cutter. Neither method worked, although an increase in water flow, partial rotation, and slow pull-up of the HQ pipe was achieved. After further working of the rods, the HQ pipe was finally released and tripped with the cutter and box end missing.

While waiting for the spare cutter to arrive, a Bowen spear was deployed to try to pull the PQ. This was unsuccessful. Throughout the rest of the day, various attempts were made to free the PQ. Backing off was also attempted, but only resulted in disconnecting close to the drill floor and not below the seabed. The PQ rods continued to be worked overnight into 20 June. At 0600 h, the spare cutter arrived, and shortly after 0900 h, the PQ had been successfully cut just inside the casing and the HQ string and cutter were returned to deck. By 0930 h, what was left of the PQ string was recovered onto deck.

To free the stuck casing, the 6⅝ inch casing cutter was modified by removing the centralizing ball bearings so that it was able to fit inside the buttress casing. The casing was successfully cut, allowing the team to recover the casing and rig down in anticipation of sailing to the next hole.

1Expedition 313 Scientists, 2010. Site M0028. In Mountain, G., Proust, J.-N., McInroy, D., Cotterill, C., and the Expedition 313 Scientists, Proc. IODP, 313: Tokyo (Integrated Ocean Drilling Program Management International, Inc.). doi:10.2204/iodp.proc.313.104.2010

2Expedition 313 Scientists' addresses.

Publication: 4 December 2010
MS 313-104